OpenStudy (oppskbeetlejuice):

A coat regularly costing $125 is put on sale for $75. What is the percent of discount? Round to the nearest percent. A. 67% B. 40% C. 50% D. 32%

OpenStudy (mathmale):

The following equation is true: (Original price) - (discount) = (sale price). Please figure out the amount of the discount here.

OpenStudy (mathmale):

Quoting you: "since you said this:The following equation is true: (Original price) - (discount) = (sale price). Please figure out the amount of the discount here. can I still answer the question? Of course you can. You don't need my permission to do that. Note that the original price and the sale price are both given. Find the discount, please.

OpenStudy (love_333):

40%

OpenStudy (betterthanvalentyn):

B

OpenStudy (mathmale):

Good. Correct. Next time, pls share your work.

OpenStudy (betterthanvalentyn):

125-75=50 discount. 125*40%=50, therefore answer is B) 40%

OpenStudy (dynamiz77):

125 - 75 = 50 discount ---> 50/125 = x/100 ---> 40%